# CLIX: Cross-Lingual Explanations of Idiomatic Expressions

### v1

@@ 1 Introduction
To explore the capabilities and limitations of modern NLP systems for generating cross-lingual explanations of idiomatic expressions, we construct a dataset composed of 628 English idioms and their explanations in English, Spanish and German. We explore performance of both pre-trained sequence-to-sequence models and Large Language Models (LLMs), and show that our task configuration is much more challenging than previous definition generation tasks. Unsurprisingly, we obtain the best performance with the model that has been exposed an immense amount of data, likely including some or all of the expressions in our dataset. We further provide some discussion on why naïve translation approaches do not suffice for this task.
@@
We release a new cross-lingual dataset containing English idiomatic expressions as well as explanations of their meanings in English, Spanish, and German.111All data and code available at: link upon publication
@@ 2 Related Work
Non-standard language takes many different forms ranging from slang to metaphor to idiom. To disambiguate some more general forms of non-standard language, Ni and Wang (2017) proposed a hybrid word-character sequence-to-sequence model that can disambiguate language found on social media. Kurfalı and Östling (2020) show that sufficiently different contextual embeddings can be found for both literal and idiomatic instances of a multi-word expression, thus allowing for disambiguation of the expressions in context. Qiang et al. (2023), on the other hand, try to directly paraphrase Chinese idioms such that they can replace the idiom using an infill strategy similar to the cloze task. These approaches are all tested in monolingual settings. The cross-lingual settings for which a solution might benefit language learners remain less explored.
@@
The cross-lingual version of idiomatic expression disambiguation has traditionally been approached as a translation task. In these types of approaches, sentences containing idiomatic expressions are mapped to equivalent sentences without such an expression, but models will often try to translate them literally word-by-word (Baziotis et al., 2023). Fadaee et al. (2018) explore sentence-level translation to generate an English–German idiom translation dataset. In their machine translation experiments, they ask their attention-based model to translate sentences which include idiomatic expressions, requiring the model to disambiguate the expressions in-context. Zaninello and Birch (2020) show that without the need for a more complicated model architecture, data augmentation and external linguistic resources can sizeably improve translations of multi-word expressions. There has additionally been use of adapters to assist in production of higher quality embeddings for idiomatic expressions with the goal of translation (Zeng and Bhat, 2022). Loss upweighting of sentences possibly containing idiomatic expressions as well as retrieval-augmented models have also been shown to improve translation capabilities, both in scenarios with and without idiomatic expressions (Liu et al., 2023). Most recently, Li et al. (2024) take a very similar prompting-based approach to us. However, the authors’ primary goal appears to be the creation of an ontology, they perform their experiments with languages we do not consider, and they additionally remain focused on a translation-based perspective. In contrast to all of the above approaches, we formulate the task more closely to definition generation, and explore the capabilities of NLP systems for this purpose in addition to applications in high-stakes scenarios such as education. We believe that this reformulation opens up interesting research avenues, such as investigating whether end-users benefit more from explanations, examples of usages, or a combination of both.
Definition Generation
@@
Since Noraset et al. (2017) introduced this task, new techniques have continued to be introduced to improve performance. Gadetsky et al. (2018) created embeddings from dictionary definitions, while taking into account the issue of polysemy by allowing their model to recognize that there could potentially be an infinite number of senses for a given word. In contrast, using a novel encoding scheme, Bevilacqua et al. (2020) are able to autoregressively generate definitions as well as score them in a discriminative manner. Zhang et al. (2023) generate cross-lingual definitions of words in English, Chinese, and French using prompt learning. They apply a contrastive learning objective to correct pre-trained models’ tendency to ignore this task and/or mix languages in the output. In this paper, we focus specifically on idiomatic expressions and their explanations, which allows us to be more flexible and we hypothesize is more challenging for NLP systems.
@@ 3.1 Task Definition
| Train | Development | Test | |
@@ 3.2 Dataset
We construct – to the best of our knowledge – the first dataset for cross-lingual explanation generation of idiomatic expressions. We build on the EPIE corpus Saxena and Paul (2020), composed of 25,206 sentences containing instances of 717 formal and static idiomatic expressions with definitions, and example usages.222Formal idiomatic expressions can undergo lexical changes, while static ones cannot. We select this dataset because it was in-part originally meant for detection of idiomatic expressions. As a result, it also contains expressions that are situationally idiomatic, a useful quality for judging explanation ability. We filter out expressions for which there are no gold explanations or in-context examples available, leaving us with 628. For each unique expression in the corpus, we include an explanation in English, Spanish, and German. We obtain non-English explanations by feeding the gold English explanations to Google Translate. For the test set, the English explanations and the automated translations are closely inspected and manually corrected by native speakers of all three languages to ensure they are entirely correct and not just the output of the automated translation. Any confusion about the meaning of a particular expression was referred to the native English speaker for review. These annotators are all computational linguists fluent in English and working in English speaking countries. In Tab. 1 we show the train/dev/test splits and the average word length of the idioms and explanations. Our splits are generated by random sample.
@@
We further enhance our idioms with thematic categories inspired by findings suggesting that classifying idioms facilitates the learning process Boers (2000). To do this, we build on the taxonomy developed by Rafatbakhsh and Ahmadi (2019), which suggests a list of 81 themes to classify English idioms (See full list with frequencies in App. C). For any idioms already assigned themes by Rafatbakhsh and Ahmadi (2019) (a little over 20%), we use the assigned theme. To tag the remainder of our dataset using this taxonomy, we prompt GPT-3.5 to provide us with the three most appropriate categories over three runs where the categories are provided in random order. We then eliminate duplicates in the resulting nine predictions and correct them by hand. The manual corrections were done by two English speaking annotators working together to resolve conflicts.
3.3 Models
In our experiments, we benchmark the performance of two sequence-to-sequence models, T5 (Raffel et al., 2020) and mT5 (Xue et al., 2021), as well as two LLMs, Llama3.1333Specifically, we use Llama3.1 8B Instruct (Dubey et al., 2024) and GPT-3.5 Turbo444Model version gpt-3.5-turbo-0125 (OpenAI, 2023).
@@ 4 Learning and Inference Strategies
In this section, we outline the different learning and inference strategies that we experiment with to generate cross-lingual explanations of idiomatic expressions. These include: whether performing the task end-to-end or decomposing it into explanation and translation, the type of contextual information used to enhance the input prompts, and the number of demonstrations used for LLM-based solutions. All prompt templates can be found in App. D
@@ 4.1 Explanation Strategies
Direct Explanation
To obtain an explanation for the benefit of a potential language learner, we fine-tune and/or prompt models for an explanation in what we assume to be the L1 of the learner. All inputs, however, are in English. Our training pairs in the fine-tuning setup consist of an input and label , where is set of tokens informing the model that its task is to produce an explanation in a target language , is an idiomatic expression, and is a gold explanation of the idiomatic expression. This model can then be formulated as:
Pipeline
In this configuration, we aim to generate our final explanation in two stages. The first stage is to generate an explanation in English. In our fine-tuning setup, we use T5 for this step. Where mT5 was explicitly designed to learn many languages, T5 was only trained to translate between a small set, and most of its training data was English. Thus, we feel T5 is more likely to perform well on this task. The model definition remains the same as in Eq. 1. However, the input pair is now and . In this case, our task sequence no longer provides any language information, and our explanation remains in the source language (in our case, English).
@@ 4.2 Contextual Enhancements
We experiment with different ways to add contextual information into the input prompt. This is done only for LLM models: Llama and GPT.
Sentence-Level Context
Because idiomatic expressions encode information that is potentially unrelated to their linguistic form, many researchers have used sentence-level context to assist with identification, disambiguation, and translation of idiomatic expressions (Fadaee et al., 2018; Kurfali and Östling, 2020; Saxena and Paul, 2020; Liu et al., 2023; Li et al., 2024). We experiment with including example sentences (one usage example per instance) containing the idiomatic expression to be explained at the end of our prompts in an effort to provide more information to the model.
Categorical Enhancements
In this setting we explore how the addition of small hints can help these models. We add to our prompt a category that an idiomatic expression might belong to. All categories we use come from the Rafatbakhsh and Ahmadi (2019) taxonomy.
| Model | Overall | Spanish | German | ||||||
| Sim | BLEU | ROUGE | Sim | BLEU | ROUGE | Sim | BLEU | ROUGE | |
| mT5 Direct | 38.06 | 0.37 | 9.94 | 36.32 | 0.34 | 10.59 | 39.80 | 0.47 | 9.34 |
| mT5 Pipeline | 37.13 | 0.09 | 4.90 | 37.13 | 0.18 | 9.71 | 37.81 | 0.29 | 7.80 |
| T5 Pipeline | 43.54 | 0.89 | 11.95 | 43.19 | 1.03 | 13.09 | 43.89 | 0.74 | 10.73 |
| Llama Direct Zero-Shot | 59.39 | 1.00 | 12.31 | 60.13 | 1.12 | 14.19 | 58.66 | 0.87 | 10.48 |
| Llama Pipeline Zero-Shot | 60.16 | 1.40 | 14.69 | 60.48 | 1.67 | 16.54 | 59.83 | 1.08 | 12.88 |
| GPT Direct Zero-Shot | 65.06 | 3.58 | 22.1 | 65.96 | 3.66 | 24.39 | 64.16 | 3.50 | 19.87 |
| GPT Pipeline Zero-Shot | 69.60 | 4.59 | 26.41 | 70.12 | 5.56 | 28.21 | 69.07 | 3.57 | 24.63 |
| GPT Direct 5-Shot | 71.15 | 4.04 | 23.23 | 71.42 | 4.22 | 25.09 | 70.87 | 3.86 | 21.34 |
| GPT Pipeline 5-Shot | 71.84 | 4.19 | 24.76 | 72.04 | 4.88 | 26.38 | 71.63 | 3.47 | 23.13 |
@@ 4.3 Few-Shot Prompting
To induce a category label, we ask the model to choose the most suitable label from our list of categories. If the induced label does not exactly match an existing label, we add it to our list and treat it like an existing label for only the instance in which we observe this idiomatic expression, and from this point forward, the procedure is the same as if we assume the category is known. Once we have a known category label (either given or acquired by induction), meaning we are given some idiomatic expression with category , we form an initial selection of idiomatic expressions with category label where and is the number of shots. If there are too few expressions to form this selection, we draw from the most similar categories until we obtain examples. Similarity between categories is calculated as the similarity between the labels by the same SBERT model used for our automated evaluation (Sec. 5). Once these examples are obtained, we take a final sample of from this set. This sampling strategy is necessitated by the small size of our dataset, where we may not have even a single example in our training set with the same category label as one in our validation or test set. The first step of selecting initial examples provides more uniformity in variance between instances where there are enough examples where the categories match and instances where there are too few.
@@ 5 Experiments
We compare results of the models described in Sec. 3.3 across a range of strategies outlined in Sec. 4. We aim to investigate what makes certain models better at understanding idiomatic expressions and what types of supplemental information are beneficial for in-context learning within this domain. We discuss the addition of contextual information outlined in Sec. 4.2. We also more closely examine methods of selecting good demonstrations in few-shot prompting scenarios when different levels of information on the newly-observed idiomatic expression vary, as described in Sec. 4.3.
Metrics
To evaluate our approaches we primarily follow Zhang et al. (2023). To determine the quality of the generated explanations, we calculate their similarity to the gold explanations in the English using a multilingual version of SBERT (Reimers and Gurevych, 2020). We use this metric as it is more likely to assess the meaning of the expressions. Finally, we include BLEU and ROUGE scores calculated with our gold translated explanations for the best models using SacreBLEU (Post, 2018) and ROUGE-L.
@@ 5.1 Results & Discussion
Our main results are described in Tab. 3. In the case of mT5 and T5, the quality of the results is generally unimpressive, demonstrated by the semantic similarity scores hovering around 40. However, we find that a pipeline strategy increased performance by more than 10%. Upon manually inspecting the outputs, in both cases the output is solely in the indicated language, but we did not see a significant difference in the fluency between the different strategies. There were, however, cases where the model would continuously repeat a key word or phrase in its response. Examples include Una persona típicamente atractiva, atractiva o atractiva and Un estado de desorientación o desorientación.
@@
In our experiments using LLMs, we see significant improvements in performance, reaching a similarity of nearly 70 with the best zero-shot strategy. This result is unsurprising, as our dataset is rather small, and probably does not include enough signal to train our T5 & mT5 models successfully. In contrast, GPT and Llama have been exposed to significantly larger amounts of data, likely including many or all of the idioms in our dataset. Consistent with the results from the sequence to sequence models, pipeline strategies seem to be most effective. While the final difference in score between the 5-shot direct and pipeline approaches is very small, we see a larger gap in the zero-shot setting. We believe that the increased gains in the direct setup are due to the direct scenario requiring more internal reasoning steps, where the demonstrations appear to help make up the difference. In contrast, the pipeline method appears to approach an upper bound in similarity scores sooner. In the case of the similar scores when using Llama in a zero-shot setting, we cannot find an exact reason. However, the fact that the pipeline approach universally performs better indicates that it is most effective.
@@
In the final results of our experiments, we found that while additional context benefitted Llama in certain cases, it did not seem to help GPT. While we are unable to perfectly pinpoint the reason it was unhelpful, we hypothesize that it is a result of the model both being more capable and the lack of a direct link between the contextual enhancement and the gold explanation. This absence is particularly noticeable in relation to experiments where the model is provided categorical information. Dev set performance can be found in Tab. 4. Lastly, we see consistently higher performance by GPT when compared to Llama across all metrics. This is likely due to our use of the – small for an LLM – 8 billion parameter version of Llama, and we do not necessarily expect to see the same trend when using a larger version.
@@
When comparing our zero and few-shot experiments, we can clearly see an increase in performance as the number of shots increases. However, we find that differing selection strategies for our examples in the few-shot scenarios do not help whatsoever. The model that performed the best used the random selection strategy. This indicates to us that either the model already understands the idiomatic expressions very well, or the relatedness of examples chosen for in-context learning are largely unimportant for this task. Results calculated on our dev set can be found in Tab. 8.
@@
| Inclusion of literal meaning (Accuracy) | • Dice with death: Participar en una actividad arriesgada o peligrosa que podría resultar potencialmente en la muerte. [ENG: Participate in an activity that could cause death.] – Makes explicit mention to death, rather than seeing it as language for danger. |

### v2

@@ 1 Introduction
To explore the capabilities and limitations of modern NLP systems for generating cross-lingual explanations of idiomatic expressions, we construct two datasets of English idioms and their explanations in English, Spanish, and German; EPIE-ME, an open dataset composed of 628 unique idioms with multi-lingual explanations built from the EPIE corpus (Saxena and Paul, 2020), and Oxford-ME, composed of 6,218 unique idioms taken from the Oxford Dictionary of Idioms 4th Edition (Ayto, 2020). We explore performance of both pre-trained sequence-to-sequence models and Large Language Models (LLMs), and show that our task configuration is much more challenging than previous definition generation tasks. We further provide some discussion on why naïve translation approaches do not suffice for this task.
@@
We release a new cross-lingual dataset containing English idiomatic expressions as well as explanations of their meanings in English, Spanish, and German.111Data and code available at: link upon publication. We are not able to publish the Oxford-ME portion due to copyright.
@@ 2 Related Work
Disambiguation of Non-Standard Language. Non-standard language takes many different forms ranging from slang to metaphor to idiom. To disambiguate some more general forms of non-standard language, Ni and Wang (2017) proposed a hybrid word-character sequence-to-sequence model that can disambiguate language found on social media. Kurfalı and Östling (2020) show that sufficiently different contextual embeddings can be found for both literal and idiomatic instances of a multi-word expression, thus allowing for disambiguation of the expressions in context. Qiang et al. (2023), on the other hand, try to directly paraphrase Chinese idioms such that they can replace the idiom using an infill strategy similar to the cloze task. These approaches are all tested in monolingual settings. The cross-lingual settings for which a solution might benefit language learners remain less explored.
@@
The cross-lingual version of idiomatic expression disambiguation has traditionally been approached as a translation task. Often, sentences containing idiomatic expressions are mapped to equivalent sentences without such an expression, but models may try to translate them word-by-word (Baziotis et al., 2023). Fadaee et al. (2018) explore sentence-level translation to generate an English–German idiom translation dataset, additionally using attention-based model to translate sentences containing idiomatic expressions, requiring in-context disambiguation. Zaninello and Birch (2020) show that data augmentation and external linguistic resources can sizably improve translations of multi-word expressions. Most recently, Li et al. (2024) take a very similar prompting-based approach to us. However, the authors’ primary goal appears to be the creation of an ontology, they perform their experiments with languages we do not consider, and they additionally remain focused on a translation-based perspective. We, however, formulate the task more closely to definition generation, and explore the capabilities of NLP systems for this purpose in addition to applications in high-stakes scenarios such as education. We believe that this reformulation opens up interesting research avenues, such as investigating whether end-users benefit more from explanations, examples usage, or a combination.
@@
Definition Generation. Since Noraset et al. (2017) introduced this task, new techniques have continued to be introduced to improve performance. Gadetsky et al. (2018) created embeddings from dictionary definitions, while taking into account the issue of polysemy by allowing their model to recognize that there could potentially be an infinite number of senses for a given word. In contrast, using a novel encoding scheme, Bevilacqua et al. (2020) are able to autoregressively generate definitions as well as score them in a discriminative manner. Zhang et al. (2023) generate cross-lingual definitions of words in English, Chinese, and French using prompt learning. They apply a contrastive learning objective to correct pre-trained models’ tendency to ignore this task and/or mix languages in the output. In this paper, we focus specifically on idiomatic expressions and their explanations, which allows us to be more flexible and we hypothesize is more challenging for NLP systems.
@@ 3.1 Task Definition
| EPIE-ME | Train | Development | Test |
@@ 3.2 Dataset
We construct – to the best of our knowledge – the first two datasets for cross-lingual explanation generation of idiomatic expressions, EPIE-ME and Oxford-ME (EPIE and Oxford with Multilingual Explanations). We build on the EPIE corpus Saxena and Paul (2020), composed of 25,206 sentences containing instances of 717 formal and static idiomatic expressions with definitions, and example usages.222Formal idiomatic expressions can undergo lexical changes, while static ones cannot. We select this dataset because it was in-part originally meant for detection of idiomatic expressions. As a result, it also contains expressions that are situationally idiomatic, a useful quality for judging explanation ability. We filter out expressions for which there are no gold explanations or in-context examples available, leaving us with 628. We additionally use the Oxford Dictionary of Idioms 4th Edition (Ayto, 2020), comprising of 6,218 English idiomatic expressions. The Oxford data only comes with sentence-level context for 1,780 of the expressions (28.6%), so we prompt Llama 3.1 to generate sentences using missing expressions in the manner found in Tab. 14. For each unique expression in the corpus, we include an explanation in English, Spanish, and German. We obtain non-English explanations by feeding the gold English explanations to Google Translate. In Tab. 1 we show the train/dev/test splits and the average word length of the idioms and explanations for both datasets. Splits are generated by random sample.
Manual corrections. For the test sets, the English explanations and the automated translations are closely inspected and manually corrected by native speakers of all three languages to ensure they are entirely correct and not just the output of the automated translation. Any confusion about the meaning of a particular expression was referred to the native English speaker for review. These annotators are all computational linguists fluent in English and working in English speaking countries.
@@
Idiom categorization. For EPIE-ME, we further enhance our idioms with thematic categories inspired by findings suggesting that classifying idioms facilitates the learning process Boers (2000). To do this, we build on the taxonomy developed by Rafatbakhsh and Ahmadi (2019), which suggests a list of 81 themes to classify English idioms (See full list with frequencies in App. C). For any idioms already assigned themes by Rafatbakhsh and Ahmadi (2019) (a little over 20%), we use the assigned theme. To tag the remainder of our dataset using this taxonomy, we prompt GPT-3.5 Turbo to provide us with the three most appropriate categories over three runs where the categories are provided in random order. We then eliminate duplicates in the resulting nine predictions and correct them by hand. The manual corrections were done by two English speaking annotators working together to resolve conflicts.
Estimating noise in non-corrected examples. Finally, we examine the amount of noise in the explanations without manual correction. We do so by calculating over our verified test set the edit distance (Levenshtein, 1965) between the original output of Google Translate and the corrected version produced by our annotators, and further normalize by length following Tashima et al. (2018). We see that the amount of correction required is generally low, the largest amount being required on average for the German EPIE-ME explanations with a normalized edit distance of 0.162. More information can be found in Appendix E.
@@ 4 Learning and Inference Strategies
In this section, we outline the different learning and inference strategies that we experiment with to generate cross-lingual explanations of idiomatic expressions. We experiment with fine-tuned and few-shot generation models. Strategies include: whether performing the task end-to-end or decomposing it into explanation and translation, the type of contextual information used to enhance the input prompts, and the number of demonstrations used for LLM-based solutions. All prompt templates can be found in App. D
@@ 4.1 Explanation Strategies
Direct Explanation. To obtain an explanation for the benefit of a potential language learner, we fine-tune and/or prompt models for an explanation in what we assume to be the L1 of the learner. All inputs, however, are in English. Our training pairs in the fine-tuning setup consist of an input and label , where is set of tokens informing the model that its task is to produce an explanation in a target language , is an idiomatic expression, and is a gold explanation of the idiomatic expression. This model can then be formulated as:
Pipeline. In this configuration, we aim to generate our final explanation in two stages. The first stage is to generate an explanation in English. In our fine-tuning setup, we use T5 for this step. Where mT5 was explicitly designed to learn many languages, T5 was only trained to translate between a small set, and most of its training data was English. Thus, we feel T5 is more likely to perform well on this task. The model definition remains the same as in Eq. 1. However, the input pair is now and . In this case, our task sequence no longer provides any language information, and our explanation remains in the source language (in our case, English).
@@ 4.2 Contextual Enhancements
We explore different ways to add contextual information into the input prompt. This is done only for LLM models: Llama and GPT.
Sentence-Level Context. Because idiomatic expressions encode information that is potentially unrelated to their linguistic form, many researchers have used sentence-level context to assist with identification, disambiguation, and translation of idiomatic expressions (Fadaee et al., 2018; Kurfali and Östling, 2020; Saxena and Paul, 2020; Liu et al., 2023; Li et al., 2024). We experiment with including example sentences (one usage example per instance) containing the idiomatic expression to be explained at the end of our prompts in an effort to provide more information to the model.
Categorical Enhancements. In this setting we explore how the addition of small hints can help these models. We add to our prompt a category that an idiomatic expression might belong to. All categories we use come from the Rafatbakhsh and Ahmadi (2019) taxonomy.
| Model | EPIE-ME | Oxford-ME | ||||
| Overall | Spanish | German | Overall | Spanish | German | |
| mT5 Direct | 38.06 | 36.32 | 39.80 | 43.21 | 43.56 | 42.86 |
| mT5 Pipeline | 37.13 | 37.13 | 37.81 | 42.58 | 42.40 | 42.77 |
| T5 Pipeline | 43.54 | 43.19 | 43.89 | 46.09 | 45.96 | 46.22 |
| Llama Direct Zero-Shot | 59.39 | 60.13 | 58.66 | 55.32 | 56.01 | 54.62 |
| Llama Pipeline Zero-Shot | 60.16 | 60.48 | 59.83 | 55.01 | 55.10 | 54.91 |
| GPT Direct Zero-Shot | 65.06 | 65.96 | 64.16 | 61.03 | 62.02 | 60.03 |
| GPT Pipeline Zero-Shot | 69.60 | 70.12 | 69.07 | 66.10 | 66.28 | 65.91 |
| GPT Direct 5-Shot | 71.15 | 71.42 | 70.87 | 66.13 | 66.9 | 65.36 |
| GPT Pipeline 5-Shot | 71.84 | 72.04 | 71.63 | 68.54 | 68.87 | 68.20 |
@@ 4.3 Few-Shot Prompting
To induce a label, we ask the model to choose the most suitable label from our list of categories. If the induced label does not exactly match an existing label, we add it to our list and treat it as an existing label for only the instance in which we observe this idiomatic expression, and from this point forward, the procedure is the same as if we assume the category is known. Once we have a known category label (either given or acquired by induction), meaning we are given some idiomatic expression with category , we form an initial selection of idiomatic expressions with category label where and is the number of shots. If there are too few expressions to form this selection, we draw from the most similar categories until we obtain examples. Similarity between categories is calculated as the similarity between the labels by the same SBERT model used for our evaluation in Sec. 5. Once these examples are obtained, we take a final sample of from this set. This sampling strategy is necessitated by the small size of our dataset, where we may not have even a single example in our training set with the same category label as one in our validation or test set. The first step of selecting initial examples helps reduce variance between instances where there are enough examples where the categories match and instances where there are too few.
@@ 5 Experiments
We compare results of several models across a range of strategies outlined in Sec. 4. We aim to investigate what makes certain models better at understanding idiomatic expressions and what types of supplemental information are beneficial for in-context learning within this domain. We discuss the addition of contextual information outlined in Sec. 4.2. We also more closely examine methods of selecting good demonstrations in few-shot prompting scenarios when levels of information for the newly-observed idiomatic expression vary, as described in Sec. 4.3.
Models & Metrics. In our experiments, we benchmark the performance of two sequence-to-sequence models, T5 (Raffel et al., 2020) and mT5 (Xue et al., 2021), as well as two LLMs, Llama3.1333Llama3.1 8B Instruct (Dubey et al., 2024) and GPT-3.5 Turbo444Model Version: gpt-3.5-turbo-0125 (OpenAI, 2023). To evaluate our approaches we primarily follow Zhang et al. (2023). To determine the quality of the generated explanations, we calculate their similarity to the gold explanations in the English using a multilingual version of SBERT (Reimers and Gurevych, 2020). We use this metric as it is more likely to assess the meaning of the expressions. Finally, in Tab. 6 we include BLEU and ROUGE scores calculated with our gold translated explanations for the best models using SacreBLEU (Post, 2018) and ROUGE-L.
@@ 5.1 Results & Discussion
Our main results are described in Tab. 3. In the case of mT5 and T5, the quality of the results is generally unimpressive, demonstrated by the semantic similarity scores hovering around 40. However, we find that a T5-based pipeline strategy increased performance by more than 10%. Upon manually inspecting the outputs, in both cases the output is solely in the indicated language, but we did not see a significant difference in the fluency between the different strategies. There were, however, cases where the model would continuously repeat a key word or phrase in its response. Examples include Una persona típicamente atractiva, atractiva o atractiva and Un estado de desorientación o desorientación. In the case of the models trained with Oxford-ME data, we only see a 10% increase in performance, indicating that models trained on the smaller EPIE-ME dataset remain competitive with those trained on 10 times the amount of data.
@@
In our experiments using LLMs, we see significant improvements in performance, reaching a similarity of nearly 70 with the best zero-shot strategy. This result is unsurprising, given that GPT and Llama have been exposed to significantly larger amounts of data, likely including many in context occurrences of the idioms in our dataset. Consistent with the results from the sequence to sequence models, pipeline strategies seem to be most effective. While the final difference in score between the 5-shot direct and pipeline approaches is very small, we see a larger gap in the zero-shot setting. We believe that the increased gains in the direct setup are due to the direct scenario requiring more internal reasoning steps, where the demonstrations appear to help make up the difference. In contrast, the pipeline method appears to approach an upper bound in similarity scores sooner. In the case of the similar scores when using Llama in a zero-shot setting, we cannot find an exact reason. However, the fact that the pipeline approach universally performs better indicates that it is most effective. For the Oxford-ME dataset, we observe that the improvement of LLMs over fine-tuned models is smaller. We attribute this to the gold explanations in the Oxford-ME dataset being about 33% shorter, penalizing the often verbose responses of the LLMs.
@@
In the final results of our experiments, we found that while additional context benefited Llama in certain cases, it did not help GPT. While we are unable to pinpoint why it was unhelpful, we hypothesize that it is a result of the model both being more capable and the lack of a direct link between the contextual enhancement and the gold explanation. This absence is particularly noticeable in relation to experiments where the model is provided categorical information. Dev set performance can be found in Tab. 4. Lastly, we see consistently higher performance by GPT when compared to Llama across all metrics. This is likely due to our use of the – small for an LLM – 8 billion parameter version of Llama, and we do not necessarily expect to see the same trend when using a larger version.
@@
When comparing our zero and few-shot experiments, we can clearly see an increase in performance as the number of shots increases. However, we find that differing selection strategies for our examples in the few-shot scenarios do not help whatsoever. The model that performed the best used the random selection strategy. This indicates to us that either the model already represents the idiomatic expressions very well, or the relatedness of examples chosen for in-context learning are largely unimportant for this task. Results calculated on our dev set can be found in Tab. 9.
@@
| Literal meaning (Accuracy) | • Dice with death: Participar en una actividad arriesgada o peligrosa que podría resultar potencialmente en la muerte. [ENG: Participate in an activity that could cause death.] – Makes explicit mention to death, rather than seeing it as language for danger. |